# Heterogenous graph neural networks for species distribution modeling

### v1

@@ 2 The case for Graph Neural Networks for Species Distributions
We represent the observed species detection records as a heterogeneous graph $\mathcal{G}(\mathcal{V}^{S},\mathcal{V}^{L},\mathcal{E}^{L2S}_{\text{Detection}},\mathcal{E}^{S2L}_{\text{Detection}})$ with distinct node sets for species $\mathcal{V}^{S}$, and locations $\mathcal{V}^{L}$. We represent detections as edges ($\mathcal{E}^{L2S}_{\text{Detection}},\mathcal{E}^{S2L}_{\text{Detection}}$) that connect species to location nodes in both directions. For location $v^{L}_{i}\in\mathcal{V}^{L}$ and species $v^{S}_{j}\in\mathcal{V}^{S}$, we include edges $e^{L2S}_{i,j}$ and $e^{S2L}_{j,i}$ in the edge set if species $j$ is detected at location $i$ 1(a)). Our learning task is a link-prediction task (Zhang & Chen, 2018) to infer whether an unseen ”detection” edge $e^{L2S}_{i^{\prime},j^{\prime}}$ should exist between location node $v^{L}_{i^{\prime}}$ and species node $v^{S}_{j^{\prime}}$. One way to compute the probability of $e^{L2S}_{i^{\prime},j^{\prime}}$ is to decode from learned representations of $v^{L}_{i^{\prime}}$ and $v^{S}_{j^{\prime}}$, where the species representation is computed by aggregating information from the location nodes with known detections through $\mathcal{E}^{L2S}_{\text{Detection}}$. More details can be found in Appendix A.2.1. The training graph may also include a separate non-detection edge set to enable message passing between species and locations without observed detections, such that separate model weights can be learned for aggregating information along these edges.
@@ 3 Application of GNNs to the NCEAS Benchmarks
(a) Example of bipartite heterogeneous graph structure where species nodes are connected to location nodes through detection edges. Solid lines represent where detection records link a species node to a location node.
(b) $\text{AUC}_{\text{ROC}}$ averaged across species per site by region and model methodology. The values of prior results were taken from the top scoring models in Valavi et al. (2022). Top result per region highlighted in blue.
For each of the six regions, the dataset provides between 11 and 13 pre-extracted environmental features, in addition to the spatial coordinate information for each data point. The goal is to predict the presence-absence status of between 20 and 54 species, depending on the region. The dataset also supplies 10,000 sampled background locations per region as part of the training data.
@@ 3.5 Results
A summary of the top $\text{AUC}_{\text{ROC}}$ models by model type, including GNNs, the baseline MLP, and prior results from Valavi et al. (2022) are displayed in Figure 1(b). Some key observations are:
@@
    Compared to prior results, GNN has better performance on 3 regions (SWI, CAN, AWT) and comparable performance on 2 regions (NZ, NSW), with a relative increase of 23.5% over prior methods in CAN. Only on the SA region does the GNN have lower performance. We suspect this is due to the SA dataset being particularly sensitive to over-fitting, which is also observed when training the baseline MLP on this region.
@@
    Among the GNN models, we observe the performance trade-off between one-way GNN and bidirectional GNN varies by region. Better understanding on the effects of edge directions and graph complexity will be explored in future work.
@@ 4 Discussion and future work
In this work, we provide a basic example of using heterogeneous GNNs for species distribution modeling through a bipartite graph as a proof-of-concept, and show that performance is on par with or exceeds other SDM methods for presence-only data. We are continuing to develop the approach further for more complex, feature-rich data that includes presence-absence data, more informative species traits and remote-sensing-derived environmental features, message-passing edges from species-to-species and location-to-location, weighted loss functions, and improved techniques for graph-based learning.

### v3

# Heterogeneous graph neural networks for species distribution modeling

@@ 2 The case for Graph Neural Networks for Species Distributions
We represent the observed species detection records as a heterogeneous graph $\mathcal{G}(\mathcal{V}^{S},\mathcal{V}^{L},\mathcal{E}^{L2S}_{\text{Detection}},\mathcal{E}^{S2L}_{\text{Detection}})$ with distinct node sets for species $\mathcal{V}^{S}$, and locations $\mathcal{V}^{L}$. We represent detections as edges ($\mathcal{E}^{L2S}_{\text{Detection}},\mathcal{E}^{S2L}_{\text{Detection}}$) that connect species to location nodes in both directions. For location $v^{L}_{i}\in\mathcal{V}^{L}$ and species $v^{S}_{j}\in\mathcal{V}^{S}$, we include edges $e^{L2S}_{i,j}$ and $e^{S2L}_{j,i}$ in the edge set if species $j$ is detected at location $i$ as shown in Figure 1. Our learning task is a link-prediction task (Zhang & Chen, 2018) to infer whether an unseen ”detection” edge $e^{L2S}_{i^{\prime},j^{\prime}}$ should exist between location node $v^{L}_{i^{\prime}}$ and species node $v^{S}_{j^{\prime}}$. One way to compute the probability of $e^{L2S}_{i^{\prime},j^{\prime}}$ is to decode from learned representations of $v^{L}_{i^{\prime}}$ and $v^{S}_{j^{\prime}}$, where the species representation is computed by aggregating information from the location nodes with known detections through $\mathcal{E}^{L2S}_{\text{Detection}}$. More details can be found in Appendix A.2.1. The training graph may also include a separate non-detection edge set to enable message passing between species and locations without observed detections, such that separate model weights can be learned for aggregating information along these edges.
Figure 1: Example of bipartite heterogeneous graph structure where species nodes are connected to location nodes through detection edges. The first graph on the left shows one way message passing from locations $v^{L}_{i}\in\mathcal{V}^{L}$ to species $v^{S}_{j}\in\mathcal{V}^{S}$ through detections $\mathcal{E}^{L2S}_{\text{Detection}}$. The middle bipartite graph includes the reversed edge set $\mathcal{E}^{S2L}_{\text{Detection}}$ that sends information from species nodes back to location nodes. The graph on the right includes message passing through (pseudo)-negative edges as a distinct edge set.
@@ 3 Application of GNNs to the NCEAS Benchmarks
For each of the six regions, the dataset provides between 11 and 13 pre-extracted environmental features, in addition to the spatial coordinate information for each data point. The goal is to predict the presence-absence status of between 20 and 54 species, depending on the region, given the presence-only detections in the training data. The characteristics of the training and test data for each region can be found in the Appendix in Table 2. The dataset also supplies 10,000 sampled background locations per region as part of the training data.
@@ 3.4 Baseline method
Figure 2: $\text{AUC}_{\text{ROC}}$ averaged across species per site by region and model methodology. The values of prior results were taken from the top scoring models in Valavi et al. (2022). Top result per region highlighted in blue.
@@ 3.5 Results
A summary of the top $\text{AUC}_{\text{ROC}}$ models by model type, including GNNs, the baseline MLP, and prior results from Valavi et al. (2022) are displayed in Figure 2. Some key observations are:
@@
    Compared to prior results, GNN has better performance on 3 regions (SWI, CAN, AWT) and comparable performance on 2 regions (NZ, NSW), with a relative increase of 23.5% over prior methods in CAN. Only on the SA region does the GNN have lower performance. We suspect this is due to the SA dataset being particularly sensitive to over-fitting, which is also observed when training the baseline MLP on this region, in addition to the relatively low number of unique locations in the test data (152 total).
@@
    Among the GNN models, we observe variation across regions in the performance trade-off between one-way GNN and bidirectional GNN message passing. Better understanding on the effects of edge directions and graph complexity will be explored in future work.
@@ 4 Discussion and future work
In this work, we provide a basic example of using heterogeneous GNNs for species distribution modeling through a bipartite graph as a proof-of-concept, and show that performance is on par with or exceeds other SDM methods for presence-only data. We are continuing to develop the approach further for more complex, feature-rich data that includes blending presence-only with presence-absence data for model training and graph structures, more informative species traits, and remote-sensing-derived environmental features. We are also developing improved graph-based learning techniques such as including message-passing through additional edges from species-to-species and location-to-location and weighted loss functions for accounting for different pseudo-negative strategies.
@@ A.4 Additional notes on benchmarking on NCEAS data
The number of species, number of location variables, unique locations with detections in the training data (excluding the additional 10,000 background locations), and unique locations in the presence-absence test data are shown in Table 2.
Region Code Region Species set Number of location variables Training Locations (with detections) Unique Test Locations
AWT Australian Wet Tropics, Queensland, Australia 20 birds; 20 vascular plants 13 3806 442
CAN Ontario, Canada 30 birds 11 5063 14571
NSW North-east New South Wales, Australia 7 bats; 8 diurnal birds; 2 nocturnal birds; 8 open-forest trees; 8 open-forest understorey vascular plants; 7 rainforest trees; 6 rainforest understory vascular plants; 8 small reptiles 13 3323 8746
NZ New Zealand 52 vascular plants 13 3088 19120
SA Continental Brazil, Ecuador, Colombia, Bolivia, and Peru, South America 30 vascular plants 11 1178 152
SWI Switzerland 30 trees 13 11429 10013
Table 2: Characteristics of the NCEAS datasets for each region